A car manufacturer claims that its car can accelerate from rest to 28 m/s in 14 seconds. Find the average acceleration and the distance the car has traveled while in the process of stopping



1
Expert's answer
2021-09-21T13:57:19-0400

By definition, the acceleration is the following:


"a = \dfrac{28m/s-0}{14s} = 2m/s^2"

where "0m\/s" represents the initial speed. The distance travelled during stopping is given by the kinematic equation (see https://www.khanacademy.org/science/physics/one-dimensional-motion/kinematic-formulas/a/what-are-the-kinematic-formulas):


"d = \dfrac{ (28m/s)^2-(0m/s)^2}{2\cdot 2m/s^2} = 196 m"

Answer. Acceleration 2 m/s^2, distance: 196 m.